OpenStudy (anonymous):

please solve \left| -1-x \right|=2

OpenStudy (anonymous):

-1-x = 2 what is the problem you have with this?

OpenStudy (anonymous):

whats the question?

OpenStudy (across):

\[\left|-x-1\right|=2,\]\[-x-1=2,\]\[x+1=2.\]

OpenStudy (across):

Solve the last two equations; those are your two answers.

OpenStudy (anonymous):

ok thanks :D

OpenStudy (anonymous):

final solution would be ?? just to be sure?

OpenStudy (anonymous):

what did you get on it?

OpenStudy (anonymous):

(3,-1)

OpenStudy (anonymous):

is this correct...?

OpenStudy (anonymous):

-3 and 1 should be what you get

OpenStudy (anonymous):

oh thx! Damn negative sign haha